OpenStudy (anonymous):

Need help! I'll fan and medal! Drag a fraction into each box to make the statements true. Answer choices may be used only once. _____ −(−712)= a negative number _____ −8/24= a positive number −34+ (_____ ) = a negative number 5/6 −1/2 −2/3

OpenStudy (anonymous):

The last one is supposed to be -3/4

OpenStudy (anonymous):

Im not sure I understand the one that says ___-8/24 because there is no way for it to be a positive -8/24 = -3 5/6 - 3 = -2 1/6 (if I am not mistaken) And that is the only positive fraction provided. Am I reading the question wrong?

OpenStudy (anonymous):

@Dramion it's supposed to be - 8/24 like subtracting

OpenStudy (anonymous):

Yes. Even so 5/6 - 8/24 cannot be subtracted.

OpenStudy (anonymous):

Technically

OpenStudy (anonymous):

So how am I supposed to answer the question?

OpenStudy (anonymous):

what I did was subtract -7/12 by 5/6 and got 17/12

OpenStudy (anonymous):

-17/12

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

I made a mistake. 5/6 −8/24= a positive number

OpenStudy (anonymous):

5/6 - 8/24 = 1/2

OpenStudy (anonymous):

I just did the math and all sets of fractions for #1 have positive outcomes and all sets of fractions have negative outcomes for #3

OpenStudy (anonymous):

Im afraid I can not help any further. Sorry :(